OpenStudy (anonymous):

Really need the answer to this, please help! If the equation is an ellipse, state the center point (h, k), the lengths of the major and minor axis, the vertices, the focal points (foci), the eccentricity and graph. 4x^2 + 9y^2 – 16x + 90y + 97 = 0

ganeshie8 (ganeshie8):

start by changing the given equation to standard form

OpenStudy (anonymous):

I know the center is (2, -5)

ganeshie8 (ganeshie8):

good :) u have the standard form already.. can u type it down for me also

OpenStudy (anonymous):

I got 4x(x-4)+9y(y+10)=-97, but got stuck

ganeshie8 (ganeshie8):

thats a very good start... lets continue from there

OpenStudy (anonymous):

I am unsure how to find the rest of the information they are wanting.

OpenStudy (anonymous):

Okay

ganeshie8 (ganeshie8):

Given equation : 4x^2 + 9y^2 – 16x + 90y + 97 = 0 as you did, first we group all x terms, and all y terms 4x^2 – 16x + 9y^2+ 90y + 97 = 0 4(x^2 - 4x) + 9(y^2+10y) = -97 4(x^2-2.2.x) + 9(y^2+2.5.y) = -97

ganeshie8 (ganeshie8):

above thing we need to change to the form : (x-h)^2/a^2 + (y-k)^2 /b^2 = 1, right ?